the area of the triangle is 10cm^2,and the base is x+2, and the height is x+1, solve for x
Answer by richwmiller(17219)

1/2*(x+2)*(x+1)=10
(x+2)*(x+1)=20
x^2+3x+2=20
x^2+3x-18=0
(x+6)*(x-3)=0
x=-6
x=3
Dismiss negative answers because we are dealing with real measurements.
x=3